OpenStudy (anonymous):

3y + 15 > –6 A. y > 3 B. y > –7 C. y > 13 D. y > –17

OpenStudy (anonymous):

first solve...3x+15=-6 3x=-6 -15=-21 x=-21/3 x=-7 then x>-7

OpenStudy (anonymous):

a madal and about three points

OpenStudy (anonymous):

and i will fan you

OpenStudy (anonymous):

@NikkiBella

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so the answer is sevan

OpenStudy (lizz123):

the answer is B

OpenStudy (lizz123):

3x+15=-6 3x=-6 -15=-21 x=-21/3 x=-7 then x>-7

OpenStudy (anonymous):

lizz123 copied lol